OpenStudy (anonymous):

the quantity of x plus 4 over 3 = 2

OpenStudy (anonymous):

(x+4)/3 = 2 x+4 = 6 x = 2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

4 over 3x + 2 over 3 = 6 ?